OpenStudy (anonymous):

How many hydrogen atoms are in 98.3 grammes of aluminium hydroxide? How many moles of Ca Cl2 does 2.41 x 10^24 formula represent.

OpenStudy (abb0t):

93.3 g Al(OH)\(_3\) \(\sf \times \frac{1 ~mol~Al(OH)_3}{mass~of~Al(OH)_3} \times \frac{6.022 \times 10^{23} atoms~Al(OH)_3}{1~mol~Al(OH)_3}\)

OpenStudy (anonymous):

is there a simpler way

OpenStudy (abb0t):

I just gave it to you. There's no simpler method.

OpenStudy (anonymous):

These types of questions are all about using the "conversion units". We are given 98.3g of aluminium hydroxide, first convert it to moles. Then we know that in 1 mole, there are 6.02x10^23 atoms/ molecules/ "whatever". Here we are looking for number of atoms, so 1 mole= 6.02x10^23 atoms

OpenStudy (anonymous):

I don't quite understand the second part of the question.

OpenStudy (anonymous):

k would be 93.3 x 6.022x10^23

OpenStudy (anonymous):

divided by the molar mass of aluminum hydroxide

OpenStudy (anonymous):

When converting from grams to moles, you have to divide the molar mass to cancel out the unit of grams.